# An almost-linear time decoding algorithm for quantum LDPC codes under circuit-level noise

### v1

@@ VII Conclusion
In this work, we introduced BP+BP+OTF as an almost linear-time decoder for QLDPC codes under circuit-level noise. The OTF post-processor removes qubits from the decoding graph until it achieves a tree-like structure, thereby increasing the likelihood that subsequent rounds of BP will converge. Additionally, we presented a novel method for mapping detector error models to sparser matrices while preserving critical information about circuit-level fault propagation. Numerical simulations demonstrate that using this mapping, the BP+BP+OTF decoder nearly matches the performance of state-of-the-art decoders such as BP+OSD when applied to families of bivariate bicycle codes under circuit-level noise.
The sparsification routine we propose extends beyond the BP+BP+OTF decoder. Our simulations suggest that performing an initial round of decoding on the full detector model, followed by a second round on the sparsified detector model, can enhance performance across various decoders, including plain BP and BP+OSD. The sparsified detector model features fewer loops, fewer columns, and a less redundant structure. Moreover, by mapping the soft information from the first BP round to the sparsified detector model, the second BP round is supplied with a non-uniform error channel, accelerating convergence. Our results show that this mapping to the sparsified detector model significantly reduces the number of required BP iterations across all the decoders we have investigated. We anticipate that other decoder families – for example BP+AC [16], BP+LSD [18], and BP+CB [17] – will also benefit from the sparsified detector error model.
In future work, we will explore the application of the BP+BP+OTF decoder to various code families, including surface codes, hypergraph product codes [6], and lifted product codes [7]. For surface codes, it would be particularly interesting to investigate whether the sparsified detector error model mapping can enhance the performance of existing decoders, such as those in [27]. For instance, this could involve implementing a BP+BP+Union-Find or BP+BP+MWPM approach.
@@
The OTF post-processor operates on the principle that QLDPC decoding can be enhanced by modifying the structure of the decoding graph. Specifically, Kruskal’s algorithm is used to identify and eliminate variable nodes that introduce problematic cycles in the Tanner graph. Similarly, the recently introduced BP plus guided decimation (BP+GD) decoder iteratively modifies the decoding graph by excluding variable nodes with the least uncertainty in their soft-information [30]. An interesting direction for future research would be to explore the combination of BP+OTF and BP+GD.
Given its low complexity, BP+BP+OTF is a promising candidate for real-time decoding of syndromes from experimental quantum computers. To this end, dedicated hardware implementations of the algorithm using FPGAs or ASICs will be necessary. Since the BP+BP+OTF algorithm uses standard methods like the well-established product-sum implementation of BP and Kruskal’s minimum spanning tree algorithm, it may be possible to construct such a decoder by combining existing commercially available chips. Consequently, development costs could potentially be lower compared to more specialised decoders such as BP+OSD [31].
@@ VIII Code availability
The code for the BP+OTF decoder can be found in the following Github repository: https://github.com/Ademartio/BPOTF. In future versions we will include a script for obtaining the transfer matrix, the two-stage BP process and the overall BP+BP+OTF decoder.
@@ IX Acknowledgements
This work was supported by the Spanish Ministry of Economy and Competitiveness through the MADDIE project (Grant No. PID2022-137099NBC44), by the Spanish Ministry of Science and Innovation through the project Few-qubit quantum hardware, algorithms and codes, on photonic and solid state systems (PLEC2021-008251), and by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda. J.R. is funded by the Engineering and Physical Sciences Research Council (Grant codes: EP/T001062/1 and EP/X026167/1).
@@ Appendix A The Union-Find data structure for the OTF decoder
In order to search an OTF in almost-linear time, we need to consider the union find data structure. Let us consider a dynamic list of a length equal to the number checks in the code to decode, where each element is a 2-element list, we shall name that list as DYNAMIC_LIST. The first element of each duple will indicate the root of the element and the second one will indicate its depth. At the beginning, each element $i$ will be its own root and its depth will be $1$, i.e. DYNAMIC_LIST$\left[i\right]=[i,1]$. Each element in the dynamic list will become a tree of a single element. Afterwards, we will consider columns in the parity check matrix in the order established by the posterior probabilities. Once we consider a column in the parity check matrix, we will study its non-trivial elements. Considering $j$ non-trivial elements, we will search at the root of each of them in the DYNAMIC_LIST. If the $j$ roots are different, introducing the edges emanating from the variable node which represents the column will not produce a loop in the Tanner graph. If that is the case, all the elements will be merged into a single tree and the root with the largest depth will become the root of the overall tree. Subsequently, the remaining roots will change their first element in the DYNAMIC_LIST to be the new root. Moreover, if there are two or more roots with largest depth, one will be chosen arbitrarily to be the new root of the merged tree and its depth will be increased by one.
After the first step, there will be elements in the DYNAMIC_LIST for which the first element will not satisfy DYNAMIC_LIST$\left[i\right]\left[0\right]=i$, but the new root that it has adopted once merging with other trees. Therefore, if we want to know its root, we will have to search the first element of the dynamic list of each element that we find until the condition DYNAMIC_LIST$\left[i\right]\left[0\right]=i$ is satisfied. Figure 3 portrays an example of the process of considering a column with three non-trivial elements, which results in the merging of three trees.
@@
Figure 3: Process of considering a column in the parity check matrix for the ordered Tanner forest. Green circles represent the elements of the DYNAMIC_LIST, which are labelled. For a violet line connecting a node on the bottom $j$ with a bottom on the top $i$, DYNAMIC_LIST$\left[j\right]\left[0\right]=i$. The top figure indicates three trees with three different roots on the top. The three red-circled nodes belong to the non-trivial elements of the column that is being considered. On the middle image, the value of the three roots of each considered node is found within the DYNAMIC_LIST. Since all roots are different, we can incorporate the column, as it does not produce a loop. We choose the root of the tree on the left as the overall root and, in the bottom figure, we merge the three trees by changing the first element in the DYNAMIC_LIST of the elements 2 and 3 to be 1. Moreover, as the tree on the left and the one on the right had the same depth, the depth of the overall tree increases by 1 to be 4.
@@ Appendix B Finding the transfer matrix by means of an exhaustive search
1$\triangleright$ Get columns of $H_{sdem}$ that at least share a non-trivial element with the detector error model column, $\mathbf{h}_{dem}^{i}$. The $map$ variable contains which elements of $H_{sdem}$ are retained so that the result for $H_{reduced}$ can be mapped back.
2:$H_{reduced}\leftarrow\emptyset$
3:$O_{reduced}\leftarrow\emptyset$
4:$map\leftarrow\emptyset$
@@
15$\triangleright$ Exhaustive search over the reduced matrix $H_{reduced}$ to decompose the detector error model column, $\mathbf{h}_{dem}^{i}$. $cols_{r}$ refers to the number of columns of $H_{reduced}=(\mathbf{h}_{reduced}^{1},\cdots,\mathbf{h}_{reduced}^{cols_{r}})$. Vectors $\mathbf{e}_{dem}^{i},\mathbf{e}_{reduced}^{m}$ refer to the error vectors associated to the $i$ and $m$ columns of their respective parity check matrix.
16:$\mathbf{a}^{i}\leftarrow\emptyset$
@@
18:     $combs\leftarrow$ GetCombinations($[1,2,\cdots,cols_{r}],j$)
@@
We aim the decomposition of each column of the detector error model into a minimum amount of sparsified components by means of the exhaustive search shown in Algorithm B. The algorithm starts by identifying the columns of $H_{sdem}$ that at least share a non-trivial component with the selected detector error model column, saved in the $H_{reduced}$ matrix. This is done to reduce the amount of elements to be considered in the exhaustive search. Once this is done, the algorithm starts to test if the detector vector in question can be expressed as a combination of the columns of the sparsified parity check matrix, starting from a single element in the sum and progressively increasing the amount of terms in the decomposition. This is done in the for loop of starting at line $17$. The GetCombinations($[1,2,\cdots,cols_{r}],j$) method creates a matrix containing all possible combinations with size $j$ of the elements of the vector $[1,2,\cdots,cols_{r}]$ as its rows. When we write $combs(k)$, we refer to getting a row of such matrix, i.e. a possible combination of columns with indexes given by such row. Therefore, the matrix obtained has size $\binom{cols_{r}}{j}\times j$, and we use those combinations to test the decomposition with $j$ elements and increase such value if no decomposition is found. We also impose that a decomposition candidate must coincide in its logical effect. Once such a decomposition is found, we generate the decomposition vector $\mathbf{a}^{i}$ with the method $ones(map(combs(k)))$ which adds ones at locations $map(combs(k))$, i.e. at the positions found but mapped back from the $H_{reduced}$ to the $H_{sdem}$ of interest. Note that since we exhaustively search from a minimum amount of columns to a bigger one, the obtained decomposition is expected to be consisted of a minimum amount of sparsified mechanisms.
@@
Algorithm B is then looped for all the columns of the detector error model in question and a transfer matrix $A_{tr}$ is found, relating the detector error model and the sparsified parity check matrix of interest. The exhaustive search approach seems to be demanding due to the fact that the combinatorial number $\binom{cols_{r}}{j}$ increases very fast with $j$. However, for all the codes considered in this article, the number of elements in the decomposition has not exceeded $3$ components and, thus, it is relatively fast. Importantly, this is done in a pre-processing stage and has not impact in the latency of the decoder. Note, also, that detector error model matrices have a periodic structure from round to round, as it can be seen in Figure 1 of [32], implying that the processing can be done for a single propagation round and then extrapolate the rest of the transfer matrix  [^3] Being strictly correct, the first round is special due to the fact that there is no propagation of previous rounds and, thus, such first round should be processed independently [32].. It is not within the scope of this work to optimize this decomposition, but to show that doing this kind of reduction to a sparsified noise model is beneficial to decode over circuit-level noise. Aiming at the best decomposition and finding better ways to search for those, e.g. using stabiliser simulators such as Stim, is considered future work.
@@ Appendix C Numerical simulations
The operational figure of merit we use to evaluate the performance of these quantum error correction schemes is the Logical Error Rate per syndrome cycle ($P_{L}$), i.e. the probability that a logical error has occurred after the recovery operation per syndrome extraction round [5].
Regarding the software implementations of the decoders used perform the numerical simulations have been: the BP+OSD implementation in the LDPC Github package [14, 34] (with slight modifications for handling circuit-level noise [5]) and our implementation for the proposed decoder can be found in the folowing Github https://github.com/Ademartio/BPOTF.
For the numerical Monte Carlo methods employed to estimate the logical error probability per syndrome cycle, $P_{L}$, we have applied the following rule of thumb to select the number of simulation runs, $N_{\mathrm{runs}}$ [4], to get the logical error rate for $d_{c}$ syndrome extraction rounds, $P_{L}(d_{c})$, as
$N_{\mathrm{runs}}=\frac{100}{P_{L}(d_{c})}.$ (7)
As explained in [4], under the assumption that the observed error events are independent, this results in a $95\%$ confidence interval of about $(0.8\hat{P_{L}}(d_{c}),1.25\hat{P_{L}}(d_{c}))$, where $\hat{P_{L}}(d_{c})$ refers to the empirically estimated value for the logical error rate. The final $P_{L}$ is just obtained by diving the numerically estimated value by the number of syndrome extraction cycles, i.e. $P_{L}=P_{L}(d_{c})/d_{c}$. [5].
@@ Appendix D Circuit-level noise model
  • Decoherence errors: data qubits are subjected to depolarizing noise before syndrome extraction round, i.e. $\{\mathrm{I,X,Y,Z}\}$, sampled independently with probabilities $p_{\mathrm{X}}=p_{\mathrm{Y}}=p_{\mathrm{Z}}=p/3$ and $p_{\mathrm{I}}=1-p$.
@@
  • State preparation: state preparations are followed by a Pauli operator which may flipped the state to its orthogonal one with probability $p$. Note that this reduces to substituting the preparation of the |$\ket{0}$ state by |$\ket{1}$ and the preparation of the $\ket{+}$ state by $\ket{-}$, each with probability $p$.
@@
  • Idle gate (memory) locations: those are followed by a Pauli operator, $\{\mathrm{I,X,Y,Z}\}$, sampled independently with probabilities $p_{\mathrm{X}}=p_{\mathrm{Y}}=p_{\mathrm{Z}}=p/3$ and $p_{\mathrm{I}}=1-p$.

### v2

@@ VII Conclusion
In this work, we introduced BP+BP+OTF as an almost linear-time decoder for QLDPC codes under circuit-level noise. The OTF post-processor removes qubits from the decoding graph until it achieves a tree-like structure, thereby increasing the likelihood that subsequent rounds of BP will converge. Additionally, we presented a novel method for mapping detector error models to sparser matrices while preserving critical information about circuit-level fault propagation. Numerical simulations demonstrate that using this mapping, the BP+BP+OTF decoder matches the performance of state-of-the-art decoders such as BP+OSD and MWPM when applied to families of bivariate bicycle and surface codes under circuit-level noise.
The sparsification routine we propose extends beyond the BP+BP+OTF decoder. Our simulations suggest that performing an initial round of decoding on the full detector model, followed by a second round on the sparsified detector model, can enhance performance across various decoders, including plain BP and BP+OSD. The sparsified detector model features fewer loops, fewer columns, and a less redundant structure. Moreover, by mapping the soft information from the first BP round to the sparsified detector model, the second BP round is supplied with a non-uniform error channel, accelerating convergence. Our results show that this mapping to the sparsified detector model can be used to match the performance of BP+OSD-0 executed over the full DEM while significantly accelerating its speed. We anticipate that other decoder families – for example BP+AC [16], BP+LSD [18], and BP+CB [17] – will also benefit from the sparsified detector error model. Furthermore, being able to operate over a detector model that resembles a phenomenological parity check matrix may make several decoding proposals to succeed at decoding circuit-level noise when those have shown promise over the simplified phenomenological noise model.
In future work, we will explore the application of the BP+BP+OTF decoder to various code families, including colour codes, hypergraph product codes [6], and lifted product codes [7].
@@
Following this line of investigation, we observed that oscillations in the a posteriori information produced by BP can be detrimental, as they may prevent the OTF step from yielding a solution consistent with the observed syndrome. To address this issue, we adopted an ensemble-based approach that enabled reliable decoding of bivariate bicycle codes. Future work could explore alternative strategies to improve the stability and reliability of the soft information. For example, automorphism ensemble BP [22], modified versions of BP [39, 23], or techniques such as iteration buffers [19, 20] could increase the likelihood of convergence.
In addition, careful optimisation of the number of BP iterations performed on both the full detector error model and the sparsified model is essential for achieving the best performance with BP+BP+OTF. In this work, we explored iteration schedules that match the decoding performance of state-of-the-art methods. Further improvements may be possible through dynamic halting strategies, where the number of iterations is adapted based on the evolution of the a posteriori information or the density of the observed syndrome.
@@
The OTF post-processor operates on the principle that QLDPC decoding can be enhanced by modifying the structure of the decoding graph. Specifically, Kruskal’s algorithm is used to identify and eliminate variable nodes that introduce problematic cycles in the Tanner graph. Similarly, the recently introduced BP+GD decoder iteratively modifies the decoding graph by excluding variable nodes with the least uncertainty in their soft-information [40]. An interesting direction for future research would be to explore the combination of BP+OTF and BP+GD.
Given its low complexity, BP+BP+OTF is a promising candidate for real-time decoding of syndromes from experimental quantum computers. To this end, dedicated hardware implementations of the algorithm using FPGAs or ASICs will be necessary. Since the BP+BP+OTF algorithm uses standard methods like the well-established minimum-sum implementation of BP and Kruskal’s minimum spanning tree algorithm, it may be possible to construct such a decoder by combining existing commercially available chips. Consequently, development costs could potentially be lower compared to more specialised decoders such as BP+OSD [41].
@@ VIII Code availability
The code for the BP+OTF decoder can be found in the following Github repository: https://github.com/Ima96/BPOTF. In future versions we will include the implementation of the ensemble BP+BP+OTF decoder.
@@ IX Acknowledgements
This work was supported by the Spanish Ministry of Economy and Competitiveness through the MADDIE project (Grant No. PID2022-137099NBC44), by the Spanish Ministry of Science and Innovation through the project Few-qubit quantum hardware, algorithms and codes, on photonic and solid state systems (PLEC2021-008251), and by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda.
J.R. is funded by an EPSRC Quantum Career Acceleration Fellowship, grant code UKRI1224. J.R. also acknowledges the support of EPSRC grants EP/T001062/1 and EP/X026167/1.
@@ Appendix A The Union-Find data structure for the OTF decoder
[…]
After the first step, there will be elements in the DYNAMIC_LIST for which the first element will not satisfy DYNAMIC_LIST[i][$\left[i\right]\left[0\right$]=i, but the new root that it has adopted once merging with other trees. Therefore, if we want to know its root, we will have to search the first element of the dynamic list of each element that we find until the condition DYNAMIC_LIST[i][$\left[i\right]\left[0\right$]=i is satisfied. Figure 5 portrays an example of the process of considering a column with three non-trivial elements, which results in the merging of three trees.
@@
[…]
@@ Appendix B Finding the transfer matrix by means of an exhaustive search
[…]
@@
[…]
@@
18:  $combs\leftarrow$ GetCombinations($1,2,\cdots,cols_{r}],j$)
@@
We aim the decomposition of each column of the detector error model into a minimum amount of sparsified components by means of the exhaustive search shown in Algorithm B. The algorithm starts by identifying the columns of $H_{sdem}$ that at least share a non-trivial component with the selected detector error model column, saved in the $H_{reduced}$ matrix. This is done to reduce the amount of elements to be considered in the exhaustive search. Once this is done, the algorithm starts to test if the detector vector in question can be expressed as a combination of the columns of the sparsified parity check matrix, starting from a single element in the sum and progressively increasing the amount of terms in the decomposition. This is done in the for loop of starting at line $17$. The GetCombinations($1,2,\cdots,cols_{r}],j$) method creates a matrix containing all possible combinations with size $j$ of the elements of the vector $1,2,\cdots,cols_{r}]$ as its rows. When we write $combs(k)$, we refer to getting a row of such matrix, i.e. a possible combination of columns with indexes given by such row. Therefore, the matrix obtained has size $\binom{cols_{r}}{j}\times j$, and we use those combinations to test the decomposition with $j$ elements and increase such value if no decomposition is found. We also impose that a decomposition candidate must coincide in its logical effect. Once such a decomposition is found, we generate the decomposition vector $\mathbf{a}^{i}$ with the method $ones(map(combs(k)))$ which adds ones at locations $map(combs(k))$, i.e. at the positions found but mapped back from the $H_{reduced}$ to the $H_{sdem}$ of interest. Note that since we exhaustively search from a minimum amount of columns to a bigger one, the obtained decomposition is expected to be consisted of a minimum amount of sparsified mechanisms.
@@
Algorithm B is then looped for all the columns of the detector error model in question and a transfer matrix $A_{tr}$ is found, relating the detector error model and the sparsified parity check matrix of interest. The exhaustive search approach seems to be demanding due to the fact that the combinatorial number $\binom{cols_{r}}{j}$ increases very fast with $j$. Importantly, this is done in a pre-processing stage and has no impact in the latency of the decoder. Note also that detector error model matrices have a periodic structure from round to round, as it can be seen in Figure 1 of [19], implying that the processing can be done for a single propagation round and then extrapolate the rest of the transfer matrix  [^4] Being strictly correct, the first round is special due to the fact that there is no propagation of previous rounds and, thus, such first round should be processed independently [19].. It is not within the scope of this work to optimize this decomposition, but to show that doing this kind of reduction to a sparsified noise model is beneficial to decode over circuit-level noise. Aiming at the best decomposition and finding better ways to search for those, e.g. using stabiliser simulators such as Stim, is considered future work.
@@ Appendix C Numerical simulations
The operational figure of merit we use to evaluate the performance of these quantum error correction schemes is the Logical Error Rate per syndrome cycle, i.e. the probability that a logical error has occurred after the recovery operation per syndrome extraction round [5].
Regarding the software implementations of the decoders used perform the numerical simulations have been: the BP+OSD implementation in the LDPC Github package [14, 26] (with slight modifications for handling circuit-level noise [5]) and our implementation for the proposed decoder can be found in the following Github https://github.com/Ima96/BPOTF.
$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$Decoding time per round (s)0.000.000.010.010.020.020.030.030.040.040.050.050.060.060.070.070.080.08Relative Frequency
(a) BP+BP+OTF at $p=0.003$.
$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$Decoding time per round (s)0.000.000.010.010.020.020.030.030.040.040.050.050.060.060.070.070.080.08Relative Frequency
(b) BP+OSD-0 at $p=0.003$.
$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$Decoding time per round (s)0.0000.0000.0050.0050.0100.0100.0150.0150.0200.0200.0250.0250.0300.0300.0350.035Relative Frequency
(c) BP+BP+OTF at $p=0.004$.
$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$Decoding time per round (s)0.0000.0000.0050.0050.0100.0100.0150.0150.0200.0200.0250.0250.0300.0300.0350.035Relative Frequency
(d) BP+OSD-0 at $p=0.004$.
Figure 6: Runtime distributions of BP+BP+OTF and BP+OSD-0 for the $[[288,12,18]]$ bivariate bicycle code. a) and c) show BP+BP+OTF at $p=0.003$ and $p=0.004$, respectively. b) and c) show BP+OSD-0 at $p=0.003$ and $p=0.004$. The data is separated into time-bins with width that grows logarithmically with decoding time.
We also use a decimation parameter to determine how aggressively the OTF is decimated from the sparsified detector model. For bivariate bicycle codes, we observed that leaving a small probability for the nodes that do not belong to the OTF improves convergence and, thus, we set said decimation parameter at 1e91e-9. For surface codes, convergence is assured over the OTF graph, which implies that the best scenario is to set the parameter to 0 and fully decimate the sparsified detector model.
Furthermore, we observe that it is important to clip the propagated values of posteriori information from the DEM to the sparsified detector model for the sake of numerical stability. In this work, we set the clipping range at $10^{-80},1.0-10^{-80}]$.
Regarding the timing analysis, we decode $20,000$ shots using both BP+OSD-0 and BP+BP+OTF (the same shots for both of them) and measure the time required to obtain a recovery. For the ensemble version of BP+BP+OTF we define the decoding time as the time required by the first decoder of the ensemble that reaches a solution because the ensemble can be run in parallel. All simulations have been run single-threaded on an Intel Core i9 based laptop.
@@ Appendix D Circuit-level noise model
[…]
@@
  • Idle locations: those are followed by a Pauli operator, $\{\mathrm{I,X,Y,Z}\}$, sampled independently with probabilities $p_{\mathrm{X}}=p_{\mathrm{Y}}=p_{\mathrm{Z}}=p/3$ and $p_{\mathrm{I}}=1-p$.
Appendix E Statistics of the timing results
In the main text, we presented the mean average time of BP+BP+OTF and BP-OSD-0 to show the speed improvement that our implementation of the decoder offers. However, the statistics of the decoding times are also interesting for elucidating the behavior of both decoders. Figure 6 shows the histograms of the decoding times obtained from $20,000$ shots for the $[[288,12,18]]$ bivariate bicycle code. We do so for the physical error rates $p=0.003$ and $p=0.004$. We select the bigger bivariate bicycle code due to the fact that it is the most demanding code.
In such figure, we can observe that BP+OSD-0 shows a clear bimodal distribution (note the peak at the right of Figures 6b and 6d) as discussed in [20]. The distributions for the BP+BP+OTF show to be more complex, with a couple of peaks. The reason for this is the fact that we used an ensemble version of the decoder to mitigate the negative effects of the oscillations of BP a posterioris. However, the spread in this case is much smaller than that shown by BP+OSD-0, indicating that the decoder is faster not only in average, but in all cases. Finally, an important thing here is that the BP implementation we used runs serially, as commented on in the main text. Thus, a fully parallelized implementation would make the distributions for BP+BP+OTF to drift to the left as the full decoder is based on BP. On the contrary, for BP+OSD-0 such an implementation would make the timing associated to BP to drift left leaving the OSD-0 peak unchanged.
[…]
(a) BP+OSD-0 over DEM vs BP+OSD-0 over sparse model.
[…]
(b) BP+OSD-0 over DEM vs BP+BP+OSD-0.
Figure 7: Performance of BP+OSD-0 when operating over the sparsified detector model. a) We compare the standard BP+OSD-0 running over the full DEM with BP+OSD-0 only over the sparsified detector model. Several orders of magnitude are lost in logical error. b) We compare the standard BP+OSD-0 running over the full detector error model with the two-stage BP and an OSD-0 post-processor, i.e. BP+BP+OSD-0. $90\%$ of BP iterations are run on the sparsified model. We observe matching accuracy. The shading highlights the region of estimated probabilities where the likelihood ratio is within a factor of $1,000$; similar to a confidence interval.
Appendix F Running BP+OSD-0 on the sparsified detector model and a two stage BP+BP+OSD-0
As stated in the main text, a natural question arising from the sparsification technique we propose is how much information is lost when doing so. In fact, this technique can also understood as compressing the detector error model. In principle, the rank of the matrices defining both models should be the same by construction, so all possible faults and detectors can appear in both models. Anyway, in order to understand this, we conduct the following simulations for the bivariate bicycle codes in [5] and compare those to the standard BP+OSD-0 (with $10,000$ maximum minimum-sum BP iterations) running over the full DEM:
  • BP+OSD-0 fully over the sparsified detector model. Here, we are oblivious to the fact that there is a DEM that has a richer information about how the faults behave at the circuit-level and directly attempt decoding over the sparsified detector model. We just map the priors using the transfer rule discussed in the text. The maximum number of minmum-sum BP iterations is set to $10,000$.
  • Leveraging the two-stage BP proposed, resulting in a BP+BP+OSD-0 decoder. We set up the BP+BP decoder discussed in the main text running $1,000$ and $9,000$ maximum minimum-sum BP iterations over the DEM and the sparsified detector model, respectively.
Figure 7 shows the obtained results. For the first case, it can easily be observed that being completely oblivious to the fact that we are compressing a bigger model results in a significant degradation in logical error rate. Interestingly, the decoder would still work and shows logical error rate reduction with code distance, but the performance is not on par with that of decoding over the full DEM. Another interesting thing that we observe by doing this relates with the mean time required for decoding. Note that, even if the sparsified detector model is significantly smaller than the DEM (recall that their sizes are around $30\%$ of the full model), we observed very similar mean decoding times for both scenarios. We consider that this comes from the fact that even if the model is smaller, the solutions to the problem will involve more columns of the sparsified model, indicating that both BP and OSD-0 will require some more time, resulting in the same thing as running the decoder over a much bigger model.
However, the second scenario resulted to be much more interesting. In Figure 7b, we can see that the performance of the BP+BP+OSD-0 decoder matches the one of BP+OSD-0. This is done by running just $10\%$ of the BP iterations over the full DEM. Furthermore, we measure that the BP+BP+OSD-0 requires $3\times$ less average decoding time than its counterpart. This comes from the fact that even if the smaller model requires solutions with higher weight, the BP iterations over the DEM and the transference to the sparsified model bias the solutions enough for faster convergence. This information that we get from the BP iterations over the DEM and their appropriate transference to the sparsified model does also explain why this decoder is able to match the accuracy of the BP+OSD-0 decoder. To sum up, the sparsified detector model can get enough information if some BP rounds are run over the full DEM. We consider that many other decoding proposals can benefit of this fact to improve their speed or to operate properly over circuit noise if they showed promise over phenomenological noise models but failed to escalate further.
Figure 8: Representation of the process of finding an ordered Tanner forest for a specific syndrome in a surface code. In a) the syndrome is presented. In b) an ordered Tanner forest implementing Kruskal’s algorithm directly in the Tanner graph can be seen. The red lines portray the variable nodes that will prevent the BP process from converging. In c) a virtual check presented on the top is introduced to the Tanner graph so as to run a Kruskal algorithm on it. In d) a Tanner forest is found which is then used in e) for finding an error recovery, also represented with a red line.
Appendix G OTF in surface codes, considerations and virtual checks
The nature of surface codes prevents the OTF algorithm from being implemented directly into the parity check matrix. This is because variable nodes in the boundaries only have a single adjacent check. To the eyes of the modified Kruskal algorithm, variable nodes connected to a single check node cannot produce loops as it would require two adjacent checks with the same root. Consequently, all variable nodes adjacent to a single check will be accepted into any Tanner forest. This will produce a heavy detriment to the task of the second BP, as some sets of variable nodes can produce trapping sets that will prevent the second BP algorithm from preventing.
Fortunately, one can circumvent this phenomenon for surface codes by including a virtual check to the Kruskal graph algorithm. This is a check only considered in Kruskal’s algorithm section that is adjacent to all variable nodes which have parity check column weight 1. Including this check allows the Kruskal algorithm to avoid trapping sets consisting of boundary variable nodes. Figure 8 portrays the inclusion of the check node when implementing Kruskal’s algorithm for surface codes. For the simulations of surface codes, we have considered two different virtual checks, one for each different independent Tanner graph.